OpenStudy (anonymous):

Please help i have exams tomorrow! List possible zeros using the rational zero theorem. Could i find these by using synthetic division? Is it the same thing? -_- thanks.

OpenStudy (anonymous):

Let me show you an example :)

OpenStudy (anonymous):

List all of the possible zeros, use synthetic division to test the possible zeros, find an actual zero, and use the actual zero to find all the zeros of x^4-5x^3+20x-16=0

OpenStudy (anonymous):

okay well my question is y=x^3-6x^2+6x-18 Thank you for being the first person to help me after i was asking everyone for like half hour! :P

OpenStudy (anonymous):

Oh Welcome !!:)

OpenStudy (anonymous):

-x^3+6 x^2-6 x+y+18 = 0

OpenStudy (anonymous):

y = x ((x-6) x+6)-18

OpenStudy (anonymous):

Hope u got some idea?

OpenStudy (anonymous):

Thanks this helps, do you think you can help me with my next one? The question is identify the number of zeros or solutions for f(x)=12x-32x^2+40-x^3 I am working on this one & the question i just asked. p.s. don't feel like you have to help me, if you r busy :)

OpenStudy (anonymous):

Not to worry :) I will help you :)

OpenStudy (anonymous):

x^2 (32 + x) + f[x] == 40 + 12 x

OpenStudy (anonymous):

f(x)+x (x (x+32)-12) = 40

OpenStudy (anonymous):

Got it?

OpenStudy (anonymous):

:( I'm still having trouble with the first question...

OpenStudy (anonymous):

Let me know your doubt?

OpenStudy (anonymous):

i can't find the right factor for the synthetic division...or maybe i made a mistake.

OpenStudy (anonymous):

Yeah, Please make a double check with your steps